Use the Laplace transform to solve the given initial-value problem. y + 4y + 29y = δ(t-a) + δ(t-3x), y(0) = 1, y(0) = 0 y(
Answer #1

\large y''+4y'+29y= \delta(t-\pi)+\delta(t-3 \pi)

\large L(y''+4y'+29y)= L(\delta(t-\pi)+\delta(t-3 \pi) )

し(N(t-37))

\large (s^2+4s+29)L(y)= L(\delta(t-\pi))+L(\delta(t-3 \pi) )+5

\large L(y)= \frac{L(\delta(t-\pi))+L(\delta(t-3 \pi) )+5}{s^2+4s+29}

y=L s2 +4s +29

21sin(5t) - 20cos(5t)

y(t) = __ e-21 (4 e6π θ(t-3 π) sin(5 t) + 4e2 π θ(t-π) sin(5 t) + 21 sin5 t)-20 cos(5 t)

21 -2t sin(5t)ecos